Preliminary version Last update made Sat Mar 20 20:55:20 EDT 2010.

FOCUSED WEEK ON QUADRATIC FORMS AND THETA FUNCTIONS

March 22 – 26, 2010

University of Florida, Gainesville, FL 32611

SCHEDULE

\underline{Location}. All lectures are in Little Hall Room 339 (The Atrium) except for that The Ramanujan Colloquium (Wednesday, 4:00pm) which will be held in FAB 103.

⋆ ⋆ ⋆

Monday, March 22, 2010

3:00 - 3:50pm Bruce Berndt, University of Illinois The Circle Problem, The Divisor Problem, and Two Bessel Series Expansions in Ramanujan’s Lost Notebook

4:05 - 4:55pm Michael Schlosser, University of Vienna, Austria Special Commutation Relations and Combinatorial Identities

⋆ ⋆ ⋆

Tuesday, March 23, 2010

9:35 - 10:25am Atul Dixit, University of llinois Transformation Formulas Associated with Integrals Involving the Riemann Xi Function

12:50 - 1:40pm William Jagy, MSRI Berkeley Ternary Quadratic Forms

4:05 - 4:55pm Steve Milne, Ohio State University Sums of Squares, Schur Functions, and Multiple basic Hypergeometric Series

7:00pm - 9:00pm DINNER at the Hilton

⋆ ⋆ ⋆

Wednesday, March 24, 2010

9:00am OUTING trip to Payne’s Prairie

3:00 - 3:30pm Bruce Berndt, University of Illinois Questions about Four Sets of Problems from Ramanujan’s Lost Notebook

4:00 - 4:55pm Kannan Soundararajan, Stanford University Quantum Unique Ergodicity and Number Theory

8:00pm - 10:00pm PARTY at the home of Krishna and Mathura Alladi

⋆ ⋆ ⋆

Thursday, March 25, 2010

9:35 - 10:25am Sun Kim, University of Illinois Proof of Ramanujan’s Second Bessel Function Identity

10:40 - 11:30am Kannan Soundararajan, Stanford University Mean Values of Multiplicative Functions and Applications

1:55 - 2:45pm Byungchan Kim, University of Illinois Asymptotic Formulas for Some Theta Functions

3:00 - 3:50pm Brandt Kronholm, SUNY at Albany Generalized Ramanujan Congruence Properties of the Restricted Partition Function p(n, m)

⋆ ⋆ ⋆

Friday, March 26, 2010

9:35 - 10:25am George Andrews, Pennsylvania State University New Aspects of False Theta Functions

10:40 - 11:40am Kannan Soundararajan, Stanford University The Distribution of Values of zeta and L-Functions

3:00 - 3:50pm Roger Baker, Bringham Young University Asymptotic Formulas for Positive Definite Quadratic Forms